SPECTROSCOPIC SIGNATURES OF ISOMERIZATION IN THE S_{1} STATE OF C_{2}H_{2}

\underline{J. H. BARABAN}, A. H. STEEVES, R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139, USA; J. F. STANTON, Institute for Theoretical Chemistry, Departments of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712; A. J. MERER, Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

Spectroscopic signatures of isomerization between the cis and trans conformers of the S_{1} state of acetylene have been studied using a reduced dimension Discrete Variable Representation (DVR) calculation. The calculated rovibrational level structure yields insight into puzzling aspects of the Ã ← X̃ band spectrum, including patterns that reveal the effects of the low-barrier cis-trans conformational change. Large anharmonicities in modes that project along the minimum energy isomerization path are discussed in the context of the isomerization barrier shape. New high-sensitivity Laser Induced Fluorescence (LIF) spectra have also been recorded in the region of the expected transition state energy in order to detect weak transitions into severely predissociated states, uncovering levels that are in accord with the DVR predictions.